### v5

# HANDI: Hand-Centric Text-and-Image Conditioned Video Generation

Abstract
Despite the recent strides in video generation, state-of-the-art methods still struggle with elements of visual detail. One particularly challenging case is the class of videos in which the intricate motion of the hand coupled with a mostly stable and otherwise distracting environment is necessary to convey the execution of some complex action and its effects. To address these challenges, we introduce a new method for video generation that focuses on hand-centric actions. Our diffusion-based method incorporates two distinct innovations. First, we propose an automatic method to generate the motion area—the region in the video in which the detailed activities occur—guided by both the visual context and the action text prompt, rather than assuming this region can be provided manually as is now commonplace. Second, we introduce a critical Hand Refinement Loss to guide the diffusion model to focus on smooth and consistent hand poses. We evaluate our method on challenging augmented datasets based on EpicKitchens and Ego4D, demonstrating significant improvements over state-of-the-art methods in terms of action clarity, especially of the hand motion in the target region, across diverse environments and actions.
Figure 1: Illustration of our proposed Hand-Centric Text-and-Image Conditioned Video Generation (HCVG). Given an image for context and an action text prompt, our method generates video frames with precisely refined hand appearance and motion, overcoming challenges like unreasonable motion in backgrounds. Unlike baselines that extend unnecessary motion to background with rough hand structure, our approach produces motion in more reasonable area, as shown in the Motion flow visualization and refined hand details.
1 Introduction
Instructional media have become the de-facto manual for everything from “how to poach an egg” videos to full laparoscopic surgery walkthroughs [61, 11]. More than half of YouTube users now search specifically for how-to content [52]. Within this vast stream, the clips most useful for both humans and robots are those that capture goal-oriented hand manipulation: fine, dexterous interactions between hands, tools, and objects performed in ordinary, cluttered environments [81, 15, 19, 62]. Such footage accelerates human skill acquisition and has proved invaluable for imitation-based policy learning [72, 27, 10].
Obtaining task-specific videos on demand remains difficult. Retrieval systems often return clips recorded in different kitchens, workshops, or lighting, breaking visual continuity with the user’s scene [2, 20]. Single-frame generators can preserve context but sacrifice temporal dynamics [28, 55]. Recent text-and-image-to-video (TI2V) models attempt to offer both context and motion [43, 79], yet they are designed for web benchmarks featuring centred subjects against clean backgrounds [4, 5, 24, 31, 48, 59, 1, 49, 3, 13, 70]; when confronted with real household footage they hallucinate background motion and render low-fidelity hands, especially when the hands occupy only a small portion of the frame (see Fig. 1). Bridging this gap is essential for practical human-assistive and robot-assistive scenarios.
Motivated by these limitations, we cast the task as hand-centric video generation (HCVG). Given (i) a single RGB scene image that shows the workspace before execution and (ii) a text command describing the desired manipulation, HCVG requires synthesising a video that (a) alters pixels only inside the motion area—the small, task-specific region encompassing the hand, tool, and target object—while leaving the surrounding context untouched; (b) produces anatomically correct, temporally smooth hand motion; and (c) realises any downstream object-state change required by the action (e.g., transforming a whole carrot into diced pieces after cut carrot; see  Fig. 1). In another word, HCVG emphasize the same input and output modality as TI2V but focuses on animating high-quality hands from challenging, unedited real-world images. Note that, unlike instruction-based TI2V, which enjoys more information (e.g., the motion area or action trajectory) from the user [6, 14, 76], HCVG requires less user burden and requires the generator itself to localise and track this region in cluttered scenes where the hand often occupies less than a quarter of the frame.
To meet these demands we introduce HANd-centric DIffusion (HANDI), a two-stage 3-D diffusion framework. Stage 1 automatically predicts a spatio-temporal motion-area (MA) mask from the image–text pair, using a lightweight video-diffusion backbone . We develop a pipeline to automatically prepare the pseudo MA masks from training videos for training. Stage 2 conditions the backbone [23, 47] on the predicted mask to generate the instructional video clip. This stage optimises a Hand-Refinement Loss for hand-pose consistency. HRL supervises by comparing the generated and groundtruth hand skeletons, across all generated frames, introducing explicit hand structure modeling among plain pixels. This two-stage MA-aware generation framework and HRL allow HANDI to render high-fidelity hands even in the presence of severe clutter.
To benchmark HCVG, we use two large-scale hand action datasets, Ego4D [19] and EpicKitchens [15], featuring various unedited real-world visual backgrounds. In addition to various standard video generation metrics, we propose two additional metrics focusing on good MA and hand appearance. We first compare with SOTA video generation methods to show HANDI’s effectiveness, demonstrating our method’s strong performance for HCVG. Then, we take a closer look at model performance along key axes: (1) the performance of samples with large MA and motion complexity defined by hand movement, (2) the ability to capture the fine-grained motion of the hands, (3) the performance within the MA versus the whole frame, and (4) compute time needed given the two-stage nature of our method. We show that our model outperforms SOTA baselines while maintaining computation efficiency.
Our key contributions are two-fold:
  1. 1.
    We identify the unique aspects of hand-centric video generation (HCVG) and propose a new method, HANDI, that is able to address these unique challenges.
  2. 2.
    We demonstrate that HANDI is able to outperform state of the art video generation from image and text-prompt methods, along all measurement axes
2 Related Work
Position of our work. HANDI unifies the desirable properties of several research threads while removing their respective limitations. Unlike image generation work [28, 55], our method produces dynamics—temporally coherent demonstrations. Unlike text-only video models (I2V), which ignore the user’s scene and therefore hallucinate context [77, 65], HANDI conditions on the real workspace image and keeps the visual context like object appearance matched. In contrast to prompt-style Text-Image-to-Video (IT2V) generation work that require masks or trajectories from the user [14, 76], HANDI does not require those scribbles and lowers user burden, enabling scalable deployment in cluttered, real-world environments where the hand occupies only a tiny fraction of the view. Finally, compared with standard TI2V systems that animate the whole frame and blur small manipulators [43, 79], we explicitly localise the motion area and refine 3-D hand pose, yielding the right amount of high-fidelity articulation.
Instructional-video corpora. Large egocentric datasets such as Assembly101 [50], YouCook2 [81], Meccano [46], HoloAssist [60], Ego4D [19] and EPIC-Kitchens [15] capture sequences of human-object interaction and drive progress in action recognition and imitation learning. Their footage is unedited: hands are small and backgrounds cluttered.
Image- and video-based visual instruction. Context-preservation is the key to providing instruction. Viewpoint and appearance mismatches increase cognitive load for humans [9] and degrade robot policies [38]. Generating static, context-aware instructional images has been explored by LEGO [28] and GenHowTo [55], but images lack temporal cues essential for manipulation. Video generation conditioned on text alone has advanced rapidly [77, 65], yet these models produce videos with mismatched context. Text-and-image-to-video (TI2V) systems such as TI2V-Zero [43] and PIA [79] add scene context but still treat the whole frame as a canvas, resulting in background motion and distorted hands. Recent instruction-based editors (e.g., Pix2Video [8], AnimateAnything [14], DragNUWA [76]) mitigate distractions by requiring user-supplied masks or keypoint trajectories, limiting scalability.
Motion-area localisation and hand fidelity. When the manipulation locus is tiny—often $<\!25\%$ of the frame—automatic localisation becomes critical. Prior work either relies on large-scale domain randomisation [16, 54] or manual markup [14, 76] and still exhibits camera drift. General-purpose video diffusers such as VideoCrafter [12] or latent transformers [44] improve realism but do not model articulated hand structure, leading to flicker and anatomical errors. Hand-specific refiners for images (HandDiffuser [41], HandRefiner [36]) do not address temporal coherence.
3 Hand-Centric Video Generation
[…]
Method Overview
We adapt SOTA video latent diffusion techniques [42, 17, 63] to be the core of our solution to the HCVG problem. Our model has two stages: stage 1 generates the motion area mask and stage 2 generates the video. However, we use the same diffusion backbone for both stages, simplifying our discussion.
Action Text Embedding Let us first explain how we represent that action text prompt $T$ to allow us to effectively condition the generation process on it. To generate the action text embedding, $E_{text}$, we first expand $T$ using a pre-trained instruction tuning module [28] to generate a rich textual description based on the evidence that this improves conditional diffusion models [28, 55, 78, 7]. After expansion, we use the CLIP text embedding [45] to compute $E_{text}\in\mathds{R}^{N\times d}$ where $N$ is the number of tokens and $d$ is the feature dimension.
[…]
For each diffusion training forward step, following standard practices [47, 34, 35], we prepare the noisy latent video $z_{t}$ for a randomly sampled $t\in{1,...,\tau}$ where $\tau$ is a predefined max step, according to standard noise-adding process [22, 44]:
$z_{t}=\sqrt{\prod_{i=1}^{t}(1-\beta_{i})}\,z_{0}+\sqrt{1-\prod_{i=1}^{t}(1-\beta_{i})}\,\epsilon,$ (1)
where $\beta_{i}$ is a predefined coefficient that controls noise intensity at step $i$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is the sampled Gaussian noise. Then, we use a 3D UNet-based Noise Predictor [14], aligned with recent video generation work [51, 17, 63, 23, 47], to predict the noise $\mathbf{\hat{\epsilon_{t}}}$ during the denoising process. The noise predictor takes the current noisy latent sample $z_{t}$ and conditions on the $E_{text}$ for prediction. Each of its blocks contains convolution layers [30] for spatial information extraction at the per frame level, temporal convolution layers [29, 57] and temporal attention layers [71] for information fusion across all timestamps at video level, and extra cross attention layers [58] to fuse information from $E_{text}$.
Inference Starting with visual context $I$, we produce $z_{context}$ from the pre-trained VAE encoder (also by Stable Diffusion v1.5 [47]). Following [14], we duplicate $z_{context}$ for $L$ times and concatenate the prior or generated MA mask, depending on stage (§3.1). Then we apply Eq. 1 at the max noise step $\tau$ to produce the initial noisy latent video $z^{\prime}_{\tau}$ as the beginning of denoising process. The denoising process during inference iterates $\tau$ steps. At each denoising step $t$ the trained stage-specific noise predictor model conditionally predicts the noise $\hat{\epsilon}_{t}$ from $z^{\prime}_{t}$, the same as training process. We apply the DPM++ solver [35, 34] to calculate the noisy latent video at previous timestamp $z^{\prime}_{t-1}$ iteratively for $\tau$ times. After $\tau$ steps of denoising, a clean latent video $z^{\prime}_{0}$ is produced. The pretrained VAE decoder [47] is applied per-frame to produce the generated motion area mask video or generated video in pixel space. We follow Dai et al. [14] to explicitly enforce the mask after the generation process, replacing the background pixels from the context image $I$.
Innovations Extending this structure, we introduce two main innovations. First (§3.1), we develop a spatially cued approach that automatically learns the motion area mask given visual context $I$ and text prompt $T$. This is critical for HCVG as the video generation needs to focus in on the pertinent area of the video while leaving the rest of it undisturbed to avoid distraction. Furthermore, it encourages the implicit learning of object state changes that frequently occur in our class of videos. Second (§3.2), we propose a novel Hand Refinement Loss tailored to HCVG. This loss greatly helps the generated video deliver higher quality critical details of action information since the hand is the key element when demonstrating actions in this scenario. We describe these two innovations in more detail next, followed by their integration into the full loss in Sec. 3.3.
3.1 Automatically Estimating the Motion Area
Hand-centric video generation must focus its effort on the spatiotemporal region that will contain the actual hand-object interaction, to avoid manipulating the visual context and adding distraction in the background. Hence, a spatiotemporal mask is needed. Although past work has supported the use of a human-supplied spatial mask in video generation [14], we focus on automatically generating the mask because it is less onerous and more flexible. The first stage of our conditional diffusion model must produce a mask video that includes any and all parts of the video that may be used in the execution of the actor given visual context $I$ and action text prompt $T$, which are used during the diffusion process.
Figure 2: Training motion area masks are automatically created from the training videos. We flood fill the convex hulls of the hand regions, and then take the set-union of these over all frames.
For the MA generation, we assume that the relevant regions of activity map to the extent of the hands in the training videos. We hence automatically annotate the motion area in a training video using off-the-shelf SOTA hand keypoint detection [37], as shown in Fig. 2.
[…]
During training of the stage 1 noise predictors, these mask videos are used as the targets. Since we use the same diffusion model structure for both stages, we also supply a prior mask during stage 1 training. This prior mask is computed as the normalized coverage over all groundtruth masks in the training set, indicating the prior distribution of the MA. During stage 2, the soft region masks outputted from stage 1, post-processed with morphology, are incorporated as part of the conditioning for the video generation. Our experiments show the high impact of the MA (§Sec. 4).
3.2 Hand Refinement Loss
It is critical yet challenging to show detailed hand appearance with consistency in action-specific motion in hand-centric video generation: the hands are typically relatively small in the pixel-space, the hands are articulated, the motion is intricate, and the background is often cluttered. We hence propose a new Hand Refinement Loss (HRL) function $\mathcal{L}_{HR}$ that improves the subtle but critical details, emphasizing the hands (see Fig. 3). Our approach is more lightweight (applied as a loss function) than previous hand generation work which require a separate model or heavy post processing [41, 36, 73, 18] .
Let us define the representation of the hand. In a given frame $l$, let $P_{l}\in\mathds{R}^{j\times 2}\;(j\in\{0,\dots J\})$ represent the hand structure, containing the 2D coordinates of the $J$ joints. In this work, we assume there are at most two hands in one frame, aligned with popular hand-centric activity datasets (e.g., EpicKitchens and Ego4D [15, 19]). We follow a commonly used hand pose skeleton topology [37] which defines 21 joints to represent structure of one hand. Therefore, $J$ is set to 42 in all our experiments.
We compute the Hand Refinement Loss between the generated video $V^{\text{gen}}\in\mathds{R}^{L\,\boxtimes\,3}$ from stage 2 and the input training video $V^{\text{train}}\in\mathds{R}^{L\,\boxtimes\,3}$, both not including the input context image. To produce $P^{\text{gen}}_{l}$ and $P^{\text{train}}_{l}$, we first apply the SOTA hand pose detector $\Upsilon$ on both videos [37]. $\Upsilon$ is aware of the temporal information in the video and configured to produce joint coordinates for left and right hands. We normalize the joints detected by $\Upsilon$ in terms of frame size to $[0,1]$. If some joints in the skeleton topology are not detected by $\Upsilon$ (i.e., the number of detected joints is less than $J$), we set the coordinates of the missing joints in either $P^{\cdot}_{l}$ as 0. Those missing joints hence do not contribute to model optimization for this frame, but the same computational effiency is maintained.
The point of adding the hand refinement loss is to encourage the generation of realistic, articulated hand pose for the specified activity. We hence freeze the hand detector $\Upsilon$ and optimize the stage 2 noise predictor to encourage better hand rendering, as follows. Given the hand pose sequences in the training and generated videos, we compute the Mean Square Error (MSE) between them as the loss value:
$\mathcal{L}_{HR}=\frac{1}{L}\sum_{l=1}^{L}\frac{1}{J}\left\|P_{l}^{\text{gen}}-P_{l}^{\text{train}}\right\|^{2}_{F}\quad.$ (2)
Note that $\mathcal{L}_{HR}$ starts to contribute densely at the latter iterations of the training when the rough hand structure in the generated video can be detected by $\Upsilon$. Clearly, computing $\mathcal{L}_{HR}$ requires the decoded video $V^{\text{gen}}$ available; we compute this video at each training iteration directly from the denoised latent video $z^{\prime}_{0}$ (see §3.3, Eq. (4)).
Figure 3: Illustrates the Hand Refinement Loss that drives the stage 2 noise predictor to focus on the fine-grained detailed of the hand pose in the context of the interaction. Bottom pose is for illustration only. Our representation has 21 joints per hand.
Figure 4: Qualitative results comparing with baselines. Our method generates videos that contain subtle hand motion that corresponds to the action description. This is because our method focuses on the effective motion region instead of cluttered background and our method is optimized with a tailored Hand Refinement Loss. The motion flow of the generated video provides more straightforward visualization and yield the same conclusion. Please see more video results in supplementary materials.
3.3 Latent- and Pixel-Space Loss Portfolio
With our innovations in automatically generating the MA mask and in the hand refinement, we require loss computation both in the latent and the pixel space at each stage.
Latent Space We define our the noise prediction loss $\mathcal{L}_{noise}$ as follows, reflecting the conditions from input video, text, and MA mask for HCVG:
$\small\mathbf{\mathcal{L}_{noise}}=\mathbb{E}_{(V,E_{text},M_{\{prior,gen\}}),\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t)\right\|^{2}_{2}\right]$ (3)
where $V$ is the training video (including the context image $I$), $E_{text}$ is the feature representation of action prompt, $M_{\{prior,gen\}}$ is the prior or generated MA mask for stage 1 or 2 respectively, $\epsilon$ is the sampled Gaussian noise and $\epsilon_{\theta}$ is the noise predicted by the noise predictor $D_{\{1,2\}}$ conditioned on the noise-added latent video $z_{t}$ with the randomly sampled noise step $t$.
Pixel Space For the pixel-level loss $\mathcal{L}_{mIoU}$ and $\mathcal{L}_{HR}$, we first decode the denoised latent video $z^{\prime}_{0}$ to the pixel space and then compute the loss between the generated video and the target video. We obtain the clean latent video $z^{\prime}_{0}$ by the following equation:
$z^{\prime}_{0}=\frac{z_{t}-\sqrt{1-\prod_{i=1}^{t}(1-\beta_{i})}\,\epsilon}{\sqrt{\prod_{i=1}^{t}(1-\beta_{i})}\,}.$ (4)
Then, we decode the actual video $V^{\text{gen}}$ from $z^{\prime}_{0}$, using the same process we described above during inference.
Stage 1 Loss We apply a mIoU loss $\mathcal{L}_{mIoU}$ in pixel space in addition to $\mathcal{L}_{noise}$ to have the final loss:
$\mathcal{L}_{stage1}=\mathcal{L}_{noise}+\alpha\mathcal{L}_{mIoU}$ (5)
where $\alpha$ is a hyperparameter tuned empirically. The Mean Intersection over Union (mIoU) loss is defined as
$\footnotesize\mathcal{L}_{\text{mIoU}}=1-\frac{1}{L}\sum_{l=1}^{L}\frac{\sum_{i,j}\left(M_{l}^{\text{gen}}\cdot M_{l}^{\text{train}}\right)}{\sum_{i,j}\left(M_{l}^{\text{gen}}+M_{l}^{\text{train}}-M_{l}^{\text{gen}}\cdot M_{l}^{\text{train}}\right)},$ (6)
where $M_{l}^{\text{gen}}$ and $M_{l}^{\text{train}}$ are the generated and training binary masks for frame $l$, summed over spatial dimensions $i$ and $j$. $L$ denotes the number frames in the output video.
Stage 2 Loss We apply the Hand Refinement Loss $\mathcal{L}_{HR}$ in addition to $\mathcal{L}_{noise}$ to have the final loss:
$\mathcal{L}_{stage2}=\mathcal{L}_{noise}+\eta\mathcal{L}_{HR}\quad,$ (7)
where the coefficient $\eta$ is a hyperparameter.
4 Experiments
Datasets We evaluate on two raw hand-action videos corpora: Epic-Kitchens-100 (EK) [15] and Ego4D [19], using LEGO’s quality-checked splits [28]. EK has 60,429/8,857 train/test samples (2.4s avg.), and Ego4D has 85,521/9,931 (1.07s avg.). These datasets reflect real-world challenges: clutter, lighting variation, and small hand regions—spanning 97/1772 verbs, 300/4336 nouns, and 45/74 kitchen or location settings in EK/Ego4D.
Benchmark Method HS-Err. $\downarrow$ VisualSim.-Frame VisualSim.-Video Consistency SemanticSim.
FID $\downarrow$ $\text{CLIP}_{GT}$uparrow FVD $\downarrow$ EgoVLP $\uparrow$ $\text{CLIP}_{Cs.}$uparrow $\text{CLIP}_{Tx.}$uparrow BLIP $\uparrow$
EpicKitchens LFDM [42] 0.01987 39.37 0.9241 129.80 0.354 0.9826 28.37 0.235
AA [14] 0.01908 5.49 0.9588 171.29 0.338 0.9843 29.97 0.295
AVDC [75] 0.01969 140.34 0.8918 81.39 0.197 0.9582 24.66 0.116
PIA [79] 0.01826 24.70 0.9446 212.88 0.361 0.9849 30.06 0.294
Open Sora [80] 0.01968 135.34 93.823 124.52 0.187 0.9573 24.46 0.186
DynamiCrafter [68] 0.01716 43.56 0.9306 175.79 0.348 0.9131 28.22 0.288
CogVideoX [74] 0.01981 127.51 0.9362 121.06 0.214 0.9677 25.13 0.176
HANDI 0.01512 5.27 0.9590 101.89 0.377 0.9896 31.14 0.298
Ego4D LFDM [42] 0.02127 50.67 0.9204 126.71 0.535 0.9821 26.93 0.221
AA [14] 0.02393 21.83 0.9647 129.60 0.642 0.9894 28.56 0.260
AVDC [75] 0.02117 144.91 0.8816 107.82 0.261 0.9722 24.17 0.155
PIA [79] 0.02393 34.62 0.9457 104.38 0.603 0.9746 29.15 0.219
Open Sora [80] 0.02142 141.90 0.8716 117.87 0.252 0.9753 24.12 0.172
DynamiCrafter [68] 0.02203 57.21 0.9386 181.24 0.336 0.9489 26.67 0.231
CogVideoX [74] 0.03079 187.20 0.8962 165.18 0.201 0.9900 28.22 0.147
HANDI 0.01939 21.51 0.9651 103.15 0.664 0.9873 28.63 0.263
Motion Intensive LFDM [42] 0.02053 56.95 0.9254 137.44 0.303 0.9825 28.39 0.210
AA [14] 0.01764 23.93 0.9591 115.14 0.368 0.9845 30.07 0.276
AVDC [75] 0.02143 148.11 0.8933 85.97 0.204 0.9579 24.60 0.102
PIA [79] 0.01940 40.97 0.9448 217.59 0.330 0.9719 30.09 0.280
HANDI 0.01663 23.79 0.9589 114.52 0.371 0.9849 31.12 0.327
Table 1: Quantitative results on EpicKitchens [15], Ego4D [19] and a Motion Intensive subset of EpicKitchens. HANDI (ours) outperforms all baselines across all metrics on at least one benchmark. VisualSim.-Frame represents the aspect of visual similarity at frame level; VisualSim.-Video represents visual similarity at video level; SemanticSim. stands for Semantic Similarity. For each column in each benchmark section, bold represents the best performance and underline stands for the second best one.
Metrics We conduct a comprehensive quantitative evaluation. For visual similarity between generated and ground truth videos, we use the image-to-image Fréchet Inception Distance (FID) [21] and $\text{CLIP}_{GT}$ score [45] as our averaged frame-level similarity, and the Fréchet Video Distance (FVD) [56] and EgoVLP score [28, 33] at the video level. For semantic similarity between the prompt and the generated video, we adopt the frame-level $\text{CLIP}_{Tx.}$ score [66] and the video level BLIP score [28, 32]. We measure motion coherency of the generated video. We apply the Consistency CLIP score $\text{CLIP}_{Cs.}$, an average CLIP similarity between consecutive frames [45]. FVD also reflects temporal coherency. In the end, to evaluate the generated hand quality, we report Hand Structure Error (HS-Err) which is defined as in Eq. (2). We report all metrics on $L$ frames in the videos not including the context image $I$.
Implementation Details For each video, we evenly sample 16 frames where the first frame is the context image $I$ and the remaining 15 frames (i.e., $L=15$) comprise the video to be denoised during training. The VAE [47], instruction tuning module [47], text encoder [47] and hand joints detection model [37] are previous work and frozen. The noise predictors are initialized with pretrained weights [14] and optimized during training. The coefficients in the loss functions are $\alpha=0.1$ and $\eta=0.1$. We train our model for 50 epochs with batch size 48 and learning rate $5\times 10^{-5}$. These hyperparameters are chosen empirically on small subsets of the training data. We use mixed precision with bf16 on 6 $\times$ H100 GPUs.
Scope No. Mask HRL HS-Err. $\downarrow$ VisualSim.-Frame VisualSim.-Video Consistency SemanticSim.
FID $\downarrow$ $\text{CLIP}_{GT}$uparrow FVD $\downarrow$ EgoVLP $\uparrow$ $\text{CLIP}_{Cs.}$uparrow $\text{CLIP}_{Tx.}$uparrow BLIP $\uparrow$
MA 1 $\times$ $\times$ 0.02067 13.516 0.9551 172.56 0.211 0.9677 27.862 0.261
2 Stage 1 $\times$ 0.01949 13.023 0.9549 122.31 0.214 0.9650 26.840 0.271
3 $\times$ 0.01771 13.440 0.9549 169.99 0.339 0.9780 27.863 0.269
4 Stage 1 0.01504 10.692 0.9563 100.52 0.390 0.9753 28.041 0.283
5 Prior 0.01701 12.530 0.9551 132.19 0.355 0.9848 27.909 0.275
6 Prior $\times$ 0.02149 11.335 0.9539 116.09 0.352 0.9714 28.011 0.262
7 GT $\times$ 0.01804 10.847 0.9547 101.53 0.367 0.9755 28.023 0.272
8 GT $\checkmark$ 0.01762 9.816 0.9536 114.44 0.372 0.9631 28.102 0.281
Full 1 $\times$ $\times$ 0.01908 5.497 0.9588 171.29 0.338 0.9843 29.970 0.295
2 Stage 1 $\times$ 0.01940 5.570 0.9589 124.42 0.197 0.9881 29.949 0.296
3 $\times$ 0.01760 5.384 0.9589 168.85 0.338 0.9889 29.929 0.294
4 Stage 1 0.01512 5.272 0.9590 101.89 0.377 0.9896 30.055 0.298
5 Prior 0.01670 5.285 0.9588 137.01 0.354 0.9869 30.045 0.297
6 Prior $\times$ 0.02146 5.613 0.9589 161.48 0.341 0.9748 30.025 0.296
7 GT $\times$ 0.01935 5.287 0.9589 102.43 0.380 0.9842 30.050 0.297
8 GT $\checkmark$ 0.01738 5.194 0.9587 113.57 0.382 0.9748 30.096 0.295
Table 2: Ablation study on the EpicKitchens benchmark [15], examining the impact of Motion Area (MA) mask generation (“Mask”) and Hand Refinement Loss (“HRL”) on model performance. The highlighted row is our full HANDI method. When Mask is “GT”, we use the ground truth mask which is supposed to be the upper bound of the performance but we still outperform it in more than half of the metrics.
4.1 Comparisons with the State-of-the-Art
Baselines  Our baseline methods include TI2V models AnimateAnything [14], PIA [79], DynamiCrafter [68], AVDC [75] and LFDM [42], as well as T2V models CogVideoX (5B) [74], and Open Sora [80]. For space, we describe only some of these here. PIA integrates temporal alignment and conditioning layers but does not explicitly target hand precision in cluttered environments. DynamiCrafter adapts a video diffusion prior for animating still images by injecting text-aligned context representations. However, its reliance on implicit motion priors makes it susceptible to generating unnatural and extended dynamics when fine-grained hand-object interactions are required. For the T2V models, CogVideoX and Open Sora focus on generating long-form, high-motion videos but lack direct image conditioning, making them ineffective for HCVG as our results show. CogVideoX employs a 3D Variational Autoencoder and an expert transformer to improve temporal consistency but suffers from flickering artifacts across frames. Since none of the baselines are originally tailored for HCVG, we fine-tune all models on the Epic-Kitchens and Ego4D benchmarks for 50 epochs (200,000 iterations, batch size 16) for fair evaluation.
Quantitative and Qualitative Results. Table 3 details our compelling quantitative comparisons. HANDI performs best across nearly all metrics and SOTA baselines, including the motion intensive samples—the subset of EK that have top 10% largest $M_{video}$3.1). In a few cases (4 values among 24), strong baselines (AVDC and AA) perform best; yet, neither shows consistent advantage across any metrics on all benchmarks. For example, although AVDC achieves the lowest FVD in EpicKitchens (the full set and motion-intensive subsets), it ranks the third for FVD on Ego4D, which is a more challenging benchmark with more diverse actions, objects, and environments (i.e., outside of the kitchen [19]). Notably, HANDI also produces the most semantically relevant videos, reflected in the highest $\text{CLIP}_{Tx.}$ and BLIP scores, which measure alignment with the text prompt. We discuss the (strong performance in the) hand structure error below in §4.2.
We show qualitative samples in Fig. 4 (and videos in supplementary materials). The key limitations of prior TI2V methods are evident, including those performing well quantitatively (PIA, AVDC, AA). Despite their image-conditioning capabilities, these baselines suffer from unwanted background artifacts or blurry/rippled scene effects due to the lack of focus on the Motion Area. For example, AVDC generates blurry videos because it is designed for robot policy learning with low-resolution video. On the other hand, the hand quality generated baselines are rough or the motion is unnatural even if not affected by the scene effects (e.g, PIA and AA). These shortcomings stem from their lack of explicit hand-motion constraints, leading to minor flickering or unnatural hand movements. Likewise, the T2V baselines (Open Sora and CogVideoX) struggle in the HCVG setting, as expected with no image context.
Fig. 4 also shows how HANDI successfully localizes motion to the hand interaction region, avoiding unnecessary background changes and global effects. Motion flow analysis further corroborates this observation, showing that HANDI concentrates motion within the targeted region while competing methods exhibit undesired movement in surrounding areas.
Computational Efficiency Analysis. Although HANDI introduces a separate stage for motion–area prediction, its overall run-time is on par with (or better than) comparable single-stage systems. On EpicKitchens, synthesising a $256\times 256$ clip on one NVIDIA H100 takes 8.6 s. For reference, the light but less expressive AnimateAnything completes in 5.7 s [14], DynamiCrafter in 9.3 s [68], PIA in 13.8 s [79], OpenSora in 44 s [80], AVDC in 66 s [75], and CogVideoX in 108 s [74]. HANDI therefore delivers its hand-fidelity gains without imposing a meaningful latency penalty.
4.2 Ablation Study
Figure 5: The visualization for ablation study showing the independent effectiveness of Motion Area generation and Hand Refinement Loss. The red boxes are not in the generated videos. They are drawn for better illustration, showing the area that reflects the effectiveness of the ablated components.
We conduct an ablation study to investigate the effect of our two new model elements—generated Motion Area (MA) and Hand Refinement Loss (HRL); shown in Tab. 2 and Fig. 5. In the ablation, we not only assess performance on the full frames (Scope is “Full”) but also assess it on the ground truth motion area $M_{video}$3.1) since this is the detailed hand-centric part of the video we focus on in HCGV (Scope is “MA”). Note that no mask (Mask is $\times$) means the motion area is set to the full frame for stage 2 generation.
Comparing rows 1 and 3 in both scopes assesses the impact of the model generated motion area. Clearly, allowing the system to generate a motion area in which to focus the stage 2 diffusion leads to higher quality video while preventing the generation of distractions and hallucinations in the background (as shown in Fig. 5). Rows 2 and 3 compare the impact of the HRL. As shown in Fig. 5 (rightmost two columns), without HRL, the generated hand could be blurry (fig. row 1 and 3) or cut off (fig. row 2) or distorted (fig. row 4). This demonstrates the capability of the HRL for detailed hand appearance improvement. Both sets of quantitative comparisons show significant improvements when the scope is set to the motion area, indicating better quality of the generated hand details, especially for FID and FVD, which capture the perceptual quality of the generated video.
5 Conclusion
We tackle the challenges in generating videos of hands performing goal-oriented actions, an understudied but important aspect of video generation for applications like human and robot learning. These scenarios demand high-fidelity hand motion, conditioned on a single image and text prompt. Our method, HANDI, builds on diffusion-based video generation with two innovations: automatic localisation of the motion area, avoiding background hallucination, and a hand refinement loss that encourages anatomically realistic and temporal consistent hand motion. Our comparisons on real-world datasets demonstrate consistent gains over prior methods while maintaining efficiency.
Appendix A BLIP Score with Different BLIP Model Size
BLIP Score [28] leverage pre-trained BLIP model [32] to evaluate how close the generated video content is to the target action text description. In the main script, we introduce the evaluation results using the large BLIP model. In this supplementary material, we provide evaluation results from the small BLIP model as well. These results yield the same conclusion as we have in the main script — our method outperforms other baseline methods.
Dataset Method BLIP_large $\uparrow$ BLIP_base $\uparrow$
EpicKitchens LFDM [42] 0.235 0.223
AA [14] 0.295 0.298
AVDC [75] 0.116 0.106
PIA [79] 0.294 0.299
HANDI 0.298 0.335
Ego4D LFDM [42] 0.221 0.219
AA [14] 0.260 0.264
AVDC [75] 0.155 0.161
PIA [79] 0.219 0.260
HANDI 0.263 0.306
Table 3: Quantitative results on EpicKitchens [15], Ego4D [19] and a Motion Intensive subset of EpicKitchens. Our method outperforms all baselines across all metrics on at least one dataset.
Appendix B Generation of Various Actions in the Same Visual Context
We explore our model’s performance on generating different actions given the same input context image. These are novel image-action pairs that have never shown in the dataset including the test set. We illustrate one example in LABEL:fig:various_prompt that crosses two existing samples in the test set. These results further indicate that our model can generate Hand-Centric videos for various actions, emphasizing hand movement and subtle fingertip motion.